### v1

# Dynamic Benchmark Construction for Evaluating Large Language Models on Real-World Codes

Abstract
As large language models (LLMs) become increasingly integrated into software development workflows, rigorously evaluating their performance on complex, real-world code generation tasks has become essential. However, existing benchmarks often suffer from data contamination and limited test rigor, constraining their ability to reveal model failures effectively. To address these, we present Code2Bench, a systematic, end-to-end pipeline for dynamically constructing robust and contamination-resistant benchmarks from real-world GitHub repositories. Specifically, Code2Bench introduces three key innovations: (1) Automated Dynamism, achieved through periodic ingestion of recent code to minimize training data contamination; (2) Scope Graph-based dependency analysis, which enables structured classification of functions into benchmark instances with controlled dependency levels (distinguishing between Self-Contained (SC) tasks for cross-language evaluation and Weakly Self-Contained (WSC) tasks involving permitted library usage); and (3) Property-Based Testing (PBT) for the automated synthesis of rigorous, high-coverage test suites to enable thorough functional verification. Using this pipeline, we construct Code2Bench-2505, the first dynamic benchmark derived from 880 recent Python projects spanning diverse domains, comprising 1,163 code generation tasks with 100% average branch coverage on ground-truth implementations. Extensive evaluation of 16 LLMs using Code2Bench-2505 reveals that models consistently struggle with SC tasks requiring complex, non-standard logic and cross-language transfer, while showing relatively stronger performance on WSC tasks in Python. Our work introduces a contamination-resistant, language-agnostic methodology for dynamic benchmark construction, offering a principled foundation for the comprehensive and realistic evaluation of LLMs on real-world software development tasks. The codes, data and results can be found at https://code2bench.github.io/.
1 Introduction
Recently, Large Language Models (LLMs) have been increasingly integrated into software development workflows [1, 2, 3], such as Github Copilot[4] and Cursor[5]. Therefore, the accurate and realistic evaluation of their performance on complex and real-world code generation tasks is essential [6, 7, 8, 9]. This necessitates robust benchmarks capable of capturing the complexities of real-world coding tasks while providing reliable and scalable evaluation.
However, existing methods individually address only subsets of these issues, and suffer from data contamination and limited test rigor. In particular, the primary challenge ❶ is the potential for training data contamination [10, 11, 12, 13], which compromises evaluation validity. Most benchmarks  [14, 15, 16] rely on static datasets, and their static nature makes them highly susceptible to contamination as LLMs are trained on large code corpora, including publicly available benchmark data. Although dynamic “live” benchmarks [17, 18, 19, 20] have emerged (e.g., using competitive programming or user interactions), they often do not extract tasks directly from the full complexity of real-world software repositories. Meanwhile, the challenge ❷, as highlighted by EvalPlus[21], a pervasive issue with many benchmarks is their insufficient test rigor, which fundamentally constrains their ability to effectively reveal subtle model failures. Most benchmarks employ a limited number of cases that are typically manually-crafted or generated using less rigorous methods [22, 23]. Even benchmarks derived from real-world code or repository contexts  [24, 25, 20, 26, 27], often lack comprehensive, automated, rigorous testing pipelines, relying instead on limited execution-based tests or manual verification. As summarized in Table 1, existing methods fall short across key dimensions, highlighting the significant limitations remaining and the necessity to design new benchmarks.
Table 1: Comparison of Code2Bench-2505 with existing code generation benchmarks.
Benchmark Source Dynamic Deps Handled Rigorous Test Multi-lang Design
HumanEval [14] Manual
MBPP [15] Manual
EvalPlus [21] Manual
LiveCodeBench [19] Contests
RepoBench [24] Project Codebases
HumanEval-X [28] Manual
BigCodeBench [23] Synthetic
DevEval [26] Project Codebases
EvoCodeBench [20] Project Codebases
Code2Bench-2505 (Ours) Project Codebases
To address these challenges, we present Code2Bench, a novel, automated, end-to-end framework for dynamically constructing rigorous, contamination-resistant benchmark instances directly from recent real-world GitHub repositories. Code2Bench uniquely integrates key properties missing in existing approaches, offering: (1) Automated Dynamism, achieved through continuously ingest code from active GitHub repositories, focusing on recent commits. This periodic update mechanism is fundamental to minimizing training data contamination risk inherent in static benchmarks and ensures the benchmark remains relevant to the evolving software development ecosystem. (2) Scope Graph-based dependency analysis, which enables structured classification of functions into instances with controlled dependency levels (distinguishing between Self-Contained (SC) tasks for cross-language evaluation and Weakly Self-Contained (WSC) tasks involving controlled usage of specific, commonly used external libraries); and (3) Property-Based Testing(PBT) for the automated synthesis of rigorous, high-coverage test suites by dynamically generating diverse inputs, which enables the discovery of subtle errors and edge cases that limited test cases might miss. Furthermore, the language-agnostic Scope Graph analysis and the broad availability of PBT libraries across common programming languages collectively enable the multi-language extensibility. of Code2Bench.
Using Code2Bench, we construct Code2Bench-2505, the first dynamic benchmark instance derived from recent real-world Python projects. Code2Bench-2505 features 1163 code generation tasks across diverse domains, with controlled dependencies (SC/WSC) and rigorous PBT-generated tests achieving 100% branch coverage on ground-truth implementations. Evaluation of 16 state-of-the-art LLMs using Code2Bench-2505 reveals that models consistently struggle with SC tasks requiring complex and non-standard logic, while showing relatively stronger performance on WSC tasks in Python. Furthermore, we investigate the nature of LLM failures, demonstrate the effectiveness of rigorous PBT tests and real-world code source in revealing nuanced model limitations that might be missed by less challenging or static benchmarks. We summarize our contributions as follows:
  • We propose Code2Bench, a novel end-to-end pipeline for dynamically constructing rigorous benchmarks from recent real-world code, directly addressing contamination, realism, dependency control, test rigor, and construction scalability.
  • Using Code2Bench, we construct the first benchmark instance Code2Bench-2505, a high-quality dataset derived from authentic code with controlled dependencies and rigorous PBT tests.
  • We provide key insights into state-of-the-art LLM performance on Code2Bench-2505, revealing cross-language inconsistencies and surprising performance trend between SC and WSC tasks.
2 The Code2Bench Framework
Figure 1: Overview of the Code2Bench Framework.
The Code2Bench framework is an automated and end-to-end pipeline meticulously designed to address the multifaceted challenges in evaluating LLMs on complex, real-world code tasks, including data contamination, insufficient rigor, dependency handling, and multi-language support. As illustrated in Figure 1, Code2Bench operates in two main stages: Candidate Filtering and Benchmark Instance Construction, which transform raw code into structured benchmark instances ready for evaluation. The design of this pipeline is underpinned by three foundational pillars:
  • Automated Dynamism: To mitigate data contamination risks, the framework continuously ingests code from active GitHub repositories, focusing on recent updates. This automated process ensures the benchmark evolves with the real-world software ecosystem, providing fresh evaluation tasks.
  • Rigorous Control: Achieving reliable evaluation requires rigorous testing and controlled task complexity. This pillar is implemented across two stages:
    • Within the Function Filtering stage, Scope Graph-based dependency analysis systematically identify and classify functions. This process enables the creation of task instances with controlled dependency levels (SC and WSC), ensuring realistic yet manageable complexity.
    • The Benchmark Instance Construction stage employs Property-Based Testing (PBT) for automated test suite generation. PBT’s ability to synthesize diverse, constraint-aware inputs based on function properties ensures rigorous, high-coverage test cases that effectively reveal subtle functional errors and edge cases. An LLM is considered to have successfully solved a problem instance if it passes all PBT-generated tests.
  • Language-Agnostic Core: The framework features a language-agnostic core, designed to ensure comprehensive multi-language extensibility. Key components like the Scope Graph analysis, based on universal program concepts, and the Property-Based Testing methodology, applicable across languages with PBT libraries, exemplify this design principle.
Crucially, the end-to-end pipeline rigorously filters and subjects function candidates to strict quality controls in test generation, ensuring that only high-quality, challenging, and well-tested instances derived from real-world code are included in the final benchmark. Given the immense scale and continuous evolution of platforms like GitHub, this rigorous process is capable of identifying and extracting a sufficient pool of such instances for benchmark construction at scale.
2.1 Candidate Filtering
The Function Filtering stage is designed to systematically extract suitable function candidates from the vast and often noisy landscape of real-world code repositories (Figure 1, left). This process is crucial for addressing challenges like data contamination and ensuring task realism and testability. It involves a multi-step pipeline employing various analysis techniques to select functions that are realistic, non-trivial, testable, and possess controllable dependencies.
Function Recall and Temporal Filtering
To obtain authentic tasks and mitigate contamination risk, we extract functions directly from actively developed GitHub codebases. We prioritize recent code by monitoring designated repositories and analyzing recent commit history. This temporal filtering helps exclude code potentially present in models’ pre-training data.
Preprocessing and Data Structuring
Following function identification, the relevant source code is preprocessed using Tree-sitter [29] to build Abstract Syntax Trees (ASTs). From these ASTs, essential relational information for each function, such as source code, parameters, type hints, and import statements, is extracted. This step standardizes the code representation and facilitates efficient subsequent analysis and filtering based on code structure and metadata.
Scope Graph based Dependency Analysis
A key requirement for benchmark tasks is having controllable dependencies to ensure testability and reduce contamination vectors from complex project structures. Using a language-agnostic Scope Graph based reference resolution technique [30], we analyze external dependencies for each function. This method identifies “unresolved references”—identifiers used but not defined locally. Based on this analysis, functions are classified: Self-Contained (SC) functions have no unresolved external references beyond language primitives, while Weakly Self-Contained (WSC) functions rely only on a predefined set of allowed, common libraries ($L_{allowed}$). Functions with dependencies outside $L_{allowed}$ (e.g., custom project modules) are discarded, ensuring selected candidates are either independent or have manageable, common dependencies crucial for reliable and standardized evaluation. Details can be found in Appendix B.
Program Analysis for Testability and Complexity
To ensure candidate functions are functionally testable and non-trivial, we apply program analysis, details of which can be found in Appendix C. We use Control Flow Graph (CFG) analysis to filter functions lacking verifiable output (e.g., no returns or only constants). Task complexity is measured via Cyclomatic Complexity [31]. Functions whose complexity does not fall within a predefined range are excluded from the analysis. To avoid duplicates, we perform function deduplication using exact function names and AST matching.
LLM-based Semantic Filtering
Structural and dependency analysis ensures technical feasibility, but may include semantically trivial functions like simple getters/setters. To ensure the final benchmark consists of challenging, relevant programming tasks, we employ an LLM-as-a-judge [32, 33] in a final filtering stage. A capable LLM assesses the semantic relevance and complexity of remaining candidates. Trivial functions identified by the LLM judge are discarded.
2.2 Benchmark Instance Construction
This stage transforms filtered function candidates into standardized benchmark instances, each accompanied by a comprehensive test case set, an executable test runner, and a precise instruction.
Test Case Generation via Property-Based Testing
A cornerstone of Code2Bench is the generation of high-quality, rigorous test cases using PBT (Figure 1, middle PBT detail). Unlike traditional example-based testing, which relies on fixed, pre-defined input-output pairs, PBT defines abstract properties that a piece of code should satisfy for a wide range of valid inputs. For code generation evaluation, the core property we test is functional equivalence: for any valid input, the output of the generated function must match the output of the ground-truth function implementation. We leverage PBT libraries (e.g., Hypothesis for Python) that allow us to define strategies for generating diverse inputs based on the function’s parameter types and constraints inferred from the ground-truth code or its specification (e.g., type hints, refined docstrings). The strategy builder process systematically explores the input space guided by these strategies and constraints. For each input $x_{i}$ generated by the PBT strategies, we execute the ground-truth function $f(x_{i})$ to deterministically obtain the expected output. This process yields a collection of $(x_{i},f(x_{i}))$ pairs, which form the rigorous test cases for the benchmark instance. Crucially, we incorporate a quality control step: only generated test suites that achieve high branch coverage (e.g., 100%) when executed on the ground-truth function are retained. We save a fixed number of high-coverage passing test cases (e.g., 500 pairs) for each problem. This generation process produces language-agnostic test cases for SC problems consisting of only basic types and library-aware test cases for WSC problems.
Testcase Runner Generation
For each benchmark instance and target evaluation program language, a Testcase Runner script is generated. The Runner’s primary role is to load PBT input-output pairs, execute the LLM-generated code, and perform differential testing [34] by comparing outputs against the ground truth. For SC tasks, the Runner is language-specific, designed to execute the language-agnostic test cases in the target language. For WSC tasks, the Runner is library-specific, incorporating utility functions to handle library-specific data types (e.g., numpy.allclose for arrays). Across all instances, a predefined recursive deep comparison function ensures accurate output validation. A dry run is performed during benchmark construction to verify the correctness of the generated Runner.
Instruction Generation
To provide clear and accurate guidance for LLM, we construct a comprehensive instruction for each benchmark instance. Inspired by the format and docstring generation process used in R2E [35], our instruction template includes a general directive to the LLM, followed by the function signature and its corresponding docstring embedded within a code snippet block. We use an LLM (distinct from the evaluated models) to refine the original function docstring or generate one if missing. To enhance robustness and mitigate potential biases from the original or LLM-generated text, inspired by prior work [23, 36], we employ back-translation using NL-Augmenter [37] to perturb the docstrings. Crucially, docstrings are tailored to task type: language-agnostic for SC tasks to support cross-language evaluation, and lib-aware for WSC tasks to precisely specify interactions with allowed libraries using relevant types.
3 The Code2Bench-2505 Benchmark
Code2Bench-2505 is the first benchmark instance constructed using Code2Bench, focusing on Python as the primary language for problem extraction while supporting cross-language evaluation for SC tasks. Dynamically sourced from high-quality open-source repositories, it ensures relevance and mitigates training data contamination risks. We sampled 880 Python projects on GitHub with at least 500 stars, targeting mature and well-maintained codebases. To ensure data lies beyond the training scope of most existing models (knowledge cutoff: August 2024) [38, 39], we extracted function candidates from commits between August 2024 and May 2025. Through Code2Bench’s multi-stage filtering pipeline, an initial pool of 1,442 candidates was refined into a final benchmark dataset of 1,163 instances. More details and examples can be found in Appendix H. The Code2Bench-2505 exhibits the following characteristics:
  • Function Category and Complexity Distribution: The dataset includes 965 Self-Contained (SC) tasks and 198 Weakly Self-Contained (WSC) tasks, reflecting varying levels of external dependency. SC tasks exhibit higher structural complexity than WSC tasks, with average cyclomatic complexity of 5.21 vs. 2.23 and average Lines of Code (LoC) of 21.25 vs. 17.25. Tasks are classified into difficulty levels (Easy, Medium, Hard) based on a capable LLM’s judgment, considering factors beyond structural metrics to reflect perceived complexity. For SC tasks, the distribution is Easy: 69, Medium: 86, Hard: 38; for WSC tasks, it is Easy: 75, Medium: 90, Hard: 33. This diverse difficulty distribution ensures comprehensive evaluation across real-world coding challenges.
  • Test Suite Quality: For each instance, we generated an average of 480 test cases for SC, 456 for WSC. These test cases achieve an average branch coverage of 100% on the ground-truth implementations, demonstrating the comprehensiveness and rigor of our test suite.
  • Problem Diversity and Dependencies: SC problems span a wide range of programming domains, including Data Processing, Mathematics, File I/O, Time Handling, Image Utilities, Logic Control, ensuring comprehensive coverage of real-world software tasks. WSC problems extend this scope by incorporating dependencies from a controlled set of 37 external libraries, such as numpy, pandas, re, and collections, reflecting realistic usage patterns in practical development scenarios.
4 Evaluation
4.1 Experimental Setup
Models Evaluated: We evaluated a diverse set of state-of-the-art Large Language Models, including both open-source and closed-source options. The complete list is provided in Table 3.
Evaluation Metric: We adopt Pass@1 [40] as our primary evaluation metric, widely used in code generation studies [19], which effectively reflects a developer’s experience with coding assistants.
Inference Settings: All models were evaluated in a zero-shot setting without example prompt-completion pairs. For Pass@1 evaluation, we used greedy decoding (temperature 0), following standard practices in deterministic code generation[23, 21, 8]. Open-source models with fewer than 32B parameters were served using vLLM [41], while larger open-source and closed-source models were accessed via official APIs.
Evaluation Process: Each model was evaluated in an automated, isolated environment. For each problem, the model received a generated instruction (including a refined docstring and function signature). The generated code was executed using language-specific Testcase Runners against PBT-generated test cases. Differential testing compared the outputs of the generated code with the ground-truth implementation for all test cases.
Table 2: Pass@1 Performance of LLMs on SC vs. WSC problems (Python implementation).
Model Name SC (%) WSC (%)
Claude-3.7-sonnet 39.9 49.5
Gemini-2.5-flash-preview 39.9 53.0
Doubao-1-5-pro-32k-250115 36.3 53.0
DeepSeek-V3 39.3 50.5
Llama-4-maverick 39.9 53.5
Qwen3-235b-a22b 37.8 50.0
Llama-4-scout 31.1 42.9
Llama-3.3-70b-Instruct 37.3 45.6
Qwen3-32b 35.2 45.5
Qwen2.5-Coder-32b-Instruct 36.2 44.4
Gemma-3-27b-it 34.2 46.0
Mistral-small-3.1-24b-Instruct 31.6 44.4
Qwen2.5-Coder-14b-Instruct 31.6 43.4
Qwen3-8b 28.5 39.9
Llama-3.2-3b-Instruct 18.1 23.2
Qwen3-1.7b 14.5 20.2
4.2 Performance Across Multiple Languages
To investigate the cross-domain implementation consistency of LLMs, we evaluate their performance on the SC subset of Code2Bench-2505 across five programming languages using the same instructions and test cases. This setup reveals how effectively LLMs translate universal task descriptions into diverse linguistic contexts.
Table 3: Pass@1 Performance of LLMs on Self-Contained (SC) problems across different languages.
Model Name Open Source Parameters Python (%) JS (%) TS (%) Java (%) Go (%)
Claude-3.7-sonnet No - 39.9 33.7 35.2 29.5 21.2
Gemini-2.5-flash-preview No - 39.9 35.2 34.1 28.5 27.5
Doubao-1-5-pro-32k-250115 No - 36.3 34.7 35.2 25.9 28.0
DeepSeek-V3 Yes 671B 39.3 34.2 32.6 25.9 23.8
Llama-4-maverick Yes 400B 39.4 34.2 30.6 22.8 11.4
Qwen3-235b-a22b Yes 235B 37.8 29.5 33.2 27.5 22.3
Llama-4-scout Yes 109B 31.1 25.4 28.0 25.4 6.2
Llama-3.3-70b-Instruct Yes 70B 37.3 27.5 27.5 24.4 15.5
Qwen3-32b Yes 32B 35.2 28.0 30.1 24.9 10.4
Qwen2.5-Coder-32b-Instruct Yes 32B 36.2 34.2 32.1 24.4 17.1
Gemma-3-27b-it Yes 27B 34.2 29.5 28.5 22.8 21.2
Mistral-small-3.1-24b-Instruct Yes 24B 31.6 30.6 23.8 22.7 13.0
Qwen2.5-Coder-14b-Instruct Yes 14B 31.6 29.5 29.5 22.8 16.6
Qwen3-8b Yes 8B 28.5 20.7 21.2 17.1 12.0
Llama-3.2-3b-Instruct Yes 3B 18.1 9.3 7.8 6.7 3.1
Qwen3-1.7b Yes 1.7B 14.5 10.4 9.3 9.3 1.6
Table 3 summarizes the Pass@1 results for each model. Performance varies significantly across languages, with Python consistently yielding the highest scores (e.g., Claude-3.7-sonnet achieving around 40% Pass@1). In contrast, models exhibit substantial drops in other languages (e.g., DeepSeek-V3 from 39.3% in Python to 23.8% in Go). These results highlight challenges in generalizing language-agnostic task descriptions, as well as language-specific hurdles such as Go’s complexity and TypeScript/Java’s strictness. Collectively, these findings underscore significant limitations in LLMs’ ability to achieve robust cross-lingual code generation on real-world codes.
4.3 Performance With Limited Dependencies
This subsection evaluates the impact of external dependencies on LLMs’ code generation performance for WSC tasks. By comparing SC and WSC tasks, we assess models’ ability to generalize to tasks involving library-specific data structures, and intricate logic on basic types. The analysis focuses on 16 models evaluated on 193 SC and 198 WSC problems in Python.
Table 2 presents the Pass@1 results for each model on SC and WSC subsets. Contrary to the common assumption that external dependencies increase task difficulty [23], our results reveal a surprising trend: models consistently achieve higher Pass@1 scores on WSC tasks compared to SC tasks in real-world scenarios. To explain this, we analyze the structural characteristics of the tasks. SC tasks have higher average cyclomatic complexity (5.21 vs. 2.23) and Lines of Code (LoC) (21.25 vs. 17.25) than WSC tasks. This suggests that the pure logical structure and control flow complexity within real-world SC functions tend to be more intricate than the core logic within real-world WSC functions. We hypothesize that LLMs, being strong pattern matchers, are relatively proficient at replicating common library interaction patterns found in WSC tasks, even if these involve complex data structures. Conversely, they struggle more with accurately reproducing the diverse, non-standard complex logic, intricate branching, and subtle edge cases often present in SC tasks, which require deeper algorithmic understanding rather than pattern recall. The detailed analysis of failure modes in Section 4.4 provides further qualitative evidence supporting this hypothesis.
4.4 Discussion and Analysis of Failure Modes
To gain deeper insights into why LLMs succeed or fail on different task types, we conducted a detailed analysis of the failure modes observed across all evaluated models. We classify outcomes into three categories: Execution Errors (code fails to run), Test Failures (code runs but produces incorrect output), and Success (code runs and passes all tests). Figure 3 presents the average distribution of these outcomes across all evaluated models for SC tasks in different programming languages and WSC tasks in Python. A key observation is the stark contrast in failure composition between SC tasks and WSC tasks. For SC tasks across all languages, the majority of failures are Test Failures, indicating that models are relatively successful at generating syntactically correct and runnable code, but frequently struggle with implementing the precise logic and handling edge cases. This pattern aligns with our hypothesis that the primary challenge in real-world SC tasks lies in their complex internal logic. In contrast, for WSC tasks in Python, the proportion of Execution Errors is significantly higher compared to SC tasks. This suggests that models are more prone to generating code that fails to run when interacting with external libraries, indicating issues with correct API usage, imports, or handling library-specific data structures. While Test Failures still occur in WSC, the increased rate of Execution Errors supports the hypothesis that handling the details of library interactions is a significant challenge. Furthermore, within SC tasks, we observe some variation across languages; for instance, Go shows a particularly high rate of Execution Errors, likely reflecting challenges related to its strictness. This analysis of failure modes provides strong evidence supporting our earlier findings regarding the differential challenges of SC and WSC tasks and offers diagnostic insights into the nature of LLM failures on real-world code. More details can be found in Appendix I.2.
4.5 Benchmarking Real-World Complexities: Code2Bench-2505 vs. EvalPlus
Figure 2: Average Performance and Failure Type Distribution across Task Types (SC vs WSC).
Figure 3: Comparing Pass@1 on Code2Bench-2505 (Python SC) and EvalPlus (HumanEval).
To demonstrate Code2Bench’s effectiveness in evaluating LLMs on real-world code complexities and revealing specific failure modes often missed by benchmarks based on simpler tasks, we compare the performance of state-of-the-art models on Code2Bench-2505 (Python SC subset) against EvalPlus’s evaluation on HumanEval [21, 14]. Both benchmarks strive for rigorous functional verification, with EvalPlus augmenting HumanEval’s test suite using mutation-based testing and Code2Bench employing Property-Based Testing (PBT) on tasks derived directly from recent real-world code. By comparing performance on these two distinct benchmarks, we can gain crucial insights into how task source (real-world dynamic vs. manually-crafted static) and test generation methodology (PBT for complex/constrained structures vs. mutation for variations) impact the revealed model capabilities and limitations. Figure 3 presents a scatter plot visualizing the Pass@1 results for a range of models evaluated on both Code2Bench-2505 (Python SC) and EvalPlus (HumanEval).
Figure 3 reveals a significant performance gap between Code2Bench-2505 and EvalPlus (HumanEval), with models consistently achieving substantially lower Pass@1 scores on Code2Bench-2505. This indicates that tasks derived from real-world code complexities in Code2Bench-2505 are generally more challenging for current LLMs compared to HumanEval’s tasks, even when both benchmarks employ rigorous testing. The scatter plot also shows distinct evaluation profiles; while a general correlation exists, models exhibit varied relative performance, suggesting the benchmarks evaluate different facets of code generation ability. This performance disparity and the differing evaluation profiles demonstrate that Code2Bench, by sourcing tasks from the dynamic and nuanced landscape of real-world software development and applying rigorous PBT tailored for complex inputs, effectively exposes LLM limitations and failure modes specific to real-world code complexities that may not be as apparent on benchmarks based on simpler, static tasks like HumanEval.
4.6 Performance by Difficulty Levels
This subsection analyzes model performance on Code2Bench-2505 tasks based on their classified difficulty levels (Easy, Medium, Hard), providing insight into how LLM capabilities scale with task complexity. We examine the average Pass@1 performance across all evaluated models SC and WSC Python tasks across these difficulty levels. The results are summarized in Table 4. More detailed results are available in Appendix I.3. Table 4 reveals a clear trend: for both SC and WSC Python
Table 4: Average Pass@1 by Difficulty in Python.
Task Type Easy(%) Medium(%) Hard(%)
SC 41.94 28.42 28.03
WSC 65.75 36.11 15.12
tasks, average Pass@1 performance decreases significantly with increasing difficulty. Easy tasks show the highest performance, followed by Medium, with Hard tasks exhibiting substantially lower Pass@1 in both categories. This confirms that Code2Bench-2505 successfully captures a wide spectrum of real-world coding challenges, from relatively straightforward problems to highly difficult ones. While WSC tasks generally show higher Pass@1 than SC tasks across Easy and Medium levels, the performance gap narrows or reverses at the Hard level, where both types of complex tasks pose significant challenges to current LLMs. This pronounced drop in performance at higher difficulty levels highlights Code2Bench’s ability to generate challenging tasks that effectively differentiate model capabilities and expose their limitations on complex real-world coding problems.
4.7 The Effectiveness of PBT-Generated Tests
To further demonstrate the rigor and discriminative power of the PBT-generated test suites in Code2Bench-2505, we analyzed the distribution of test case pass rates. While Pass@1 (100% passing) is our primary metric, examining the percentage of problems where models achieved a test case pass rate $\geq 98\%$ (indicating $\leq 10$ failures out of $\approx 500$ tests) provides crucial insight into the capture of subtle errors. Table 5 presents this percentage for four representative models across various task types (SC tasks in Python, Go, Java, TS, JS, and WSC tasks in Python). The metric is computed as the count of problems passing $\geq 98\%$ of tests ($\leq 2\%$ failed) divided by the total number of distinct problems in each respective category.
Table 5: Percentage of Problems with Test Case Pass Rate $\geq$ 98% for Representative Models (%).
Model Name SC: Python SC: Go SC: Java SC: TS SC: JS WSC: Python
Claude-3.7-sonnet 13.5 14.7 15.6 10.9 18.6 8.8
Gemini-2.5-flash-preview 11.9 11.8 9.5 10.4 19.1 9.8
DeepSeek-V3 11.9 16.1 16.5 10.9 20.9 7.8
Qwen3-235b-a22b 12.9 10.8 10.4 10.4 19.5 7.3
As shown in Table 5, even for these capable models, the percentage of problems with a test case pass rate of 98% or higher generally ranging from approximately 7% to 21%. This highlights that models frequently produce code that is nearly correct but fails on a small number of test cases. If these specific, subtle-error-revealing test cases were absent, these problems would be incorrectly marked as perfectly solved. This underscores the effectiveness of our comprehensive PBT-generated test suites in capturing nuanced failures on complex, real-world code, providing a more reliable assessment than possible with less thorough testing. More details can be found in Appendix I.4.
5 Related Work
Code Generation Benchmarks
Evaluating Large Language Models (LLMs) on code generation tasks is an active research area, with numerous benchmarks proposed. Early benchmarks[42, 43] like HumanEval [14], MBPP [15], and APPS [16] provide static collections of isolated code snippets with tests, proving valuable for initial model development but facing limitations regarding dataset contamination [10, 11, 12, 13] and the lack of real-world context and dependencies. Efforts like EvalPlus [21] enhance static benchmarks with more robust tests via mutation. However, these may not fully represent real-world code complexities or provide automated construction from code repositories. Multilingual benchmarks[44] like HumanEval-X [28] and Aider polyglot benchmark [45] evaluate cross-language abilities but are at risk of data leakage. Benchmarks focusing on repository-level context or dependencies[46, 20, 47, 48] include RepoBench [24], CrossCodeEval [25], R2E [35], DevEval [26], BigCodeBench [23], and CODEAGENT [27]. WebBench [49] introduces sequential, real-world web development tasks. While these capture aspects of real-world interaction, they often lack sufficient testing. Code2Bench stands out by offering an end-to-end pipeline for dynamically generating rigorous benchmark instances from recent real-world GitHub repositories.
Data Leakage and Live Benchmarks
A key limitation of static benchmarks is the risk of test set contamination, where models may have been trained on the same data used for evaluation[10, 11, 12, 13]. This has motivated the development of "live" benchmarks. DynaBench [50] identified these challenges and advocated for continuously evolving benchmarks. Chatbot Arena [17] provides a platform for dynamic evaluation based on user interactions. LiveBench [18] sources new data from specific, non-code domains like mathematics and news, while LiveCodeBench [19] collects recent problems from competitive programming platforms. Other methods leverage LLMs for task mutation to generate new problems, such as EvoEval [51]. Evocodebench [20] introduces a periodically updated benchmark to mitigate leakage, but currently focuses on Python and may lack rigorous test. Arena-Hard-Auto [52] filters crowdsourced prompts into benchmarks. While existing initiatives have made progress in addressing contamination and enabling dynamic evaluation, they often rely on specific data sources, focus on evaluation platforms, or lack a systematic, automated pipeline for constructing high-quality benchmarks from real-world code at scale. Code2Bench bridges this gap by providing a novel, automated, end-to-end pipeline that dynamically extracts, filters, and constructs rigorous benchmark instances from recent activity in real-world GitHub repositories.
6 Conclusion & Future Work
We introduced Code2Bench, a novel automated framework for dynamically constructing rigorous, contamination-resistant benchmarks from real-world code. By integrating dynamism, systematic dependency control, and Property-Based Testing, Code2Bench addresses limitations of existing benchmarks and enables reliable evaluation. Our benchmark, Code2Bench-2505, revealed that state-of-the-art LLMs struggle with complex logic (SC tasks) and cross-language generalization but perform better on common library interactions (WSC tasks in Python). These results highlight specific LLM weaknesses and demonstrate the value of Code2Bench in providing deeper insights into realistic coding challenges.
Despite these advances, our work has limitations in evaluation scope and task coverage, discussed in detail in Appendix J. Future work will focus on extending Code2Bench’s capabilities. Key directions include expanding language support by leveraging the framework’s language-agnostic core and enabling domain-specific benchmark construction. The language-agnostic Scope Graph-based dependency analysis , central to our framework, is inherently extensible. By defining Scope Graph rules for domain-specific libraries and APIs (e.g., system calls in systems programming or specialized data science libraries), Code2Bench can automatically generate tasks tailored to real-world coding challenges in these domains.
Appendix A Preprocessing and Data Structuring
This appendix details the preprocessing and data structuring steps performed in the Function Filtering pipeline (Section 2) to transform raw source code into a standardized representation suitable for subsequent analysis.
A.1 Source Code Parsing via Abstract Syntax Trees (ASTs)
Following the initial identification of candidate functions from the source code repositories, the raw text code of these functions and their surrounding context is processed using Tree-sitter [29]. Tree-sitter is a parser generator tool that produces concrete syntax tree parsers for various programming languages. Unlike traditional compilers that focus on semantic analysis, Tree-sitter is specifically designed for source code analysis tools, providing robust, incremental parsing capabilities and generating detailed, well-structured Concrete Syntax Trees (CSTs), which are closely related to Abstract Syntax Trees (ASTs).
For each identified function, the corresponding source code snippet is parsed into its AST representation. The AST is a tree structure that represents the abstract syntactic structure of source code written in a programming language. Each node in the tree denotes a construct occurring in the source code (e.g., function definition, variable declaration, expression, statement). Parsing the raw code into an AST is crucial as it:
  • Standardizes the code representation, abstracting away syntactic variations (e.g., whitespace, comments) and providing a consistent structure regardless of the original formatting.
  • Exposes the hierarchical and relational structure of the code, making it amenable to systematic program analysis techniques.
A.2 Extraction of Relational Information
From the generated ASTs, we extract essential relational information and metadata for each candidate function. This extracted information is crucial for the subsequent dependency analysis, program analysis, semantic filtering, and benchmark instance construction stages. Key information extracted includes:
  • Function Signature: The function name, parameters, and their declared types or type hints (if available). This information is directly used for generating the benchmark instruction.
  • Source Code Snippet: The exact lines of code corresponding to the function definition, serving as the Ground Truth implementation and the basis for program analysis.
  • Import Statements: Identification of modules or names imported within the function’s scope or in its surrounding file. This is vital for understanding potential external dependencies.
  • Metadata: Information such as the original file path and commit hash, aiding in traceability and ensuring the recency of the code source.
This structured extraction process transforms the raw, unstructured code into a queryable and analyzable format, laying the foundation for the automated benchmark construction pipeline.
Appendix B Scope Graph Based Dependency Analysis
This appendix provides a formalized technical explanation of the Scope Graph-based dependency analysis employed in the Function Filtering stage of the Code2Bench. This analysis is fundamental to identifying and controlling external dependencies of candidate functions, enabling the classification of tasks into Self-Contained (SC) and Weakly Self-Contained (WSC) categories crucial for rigorous and standardized evaluation.
B.1 Scope Graph Model
After parsing the source code of a project into Abstract Syntax Trees (ASTs) using Tree-sitter, we construct a Scope Graph $G=(V,E)$. The vertex set $V$ includes:
  • Scope nodes $S\subseteq V$, representing hierarchical scopes like modules, classes, functions, or blocks. Each scope $s\in S$ represents a region of code where identifiers are defined and resolved.
  • Definition nodes $D\subseteq V$, representing the points where identifiers are defined (e.g., variable declarations, function definitions). Each definition $d\in D$ corresponds to a specific identifier name $id(d)$.
  • Reference nodes $R\subseteq V$, representing the points where identifiers are used (referenced) within the code. Each reference $r\in R$ corresponds to a specific identifier name $id(r)$.
The edge set $E$ includes:
  • Scope hierarchy edges $E_{scope}\subseteq S\times S$, representing nested scopes (e.g., a function scope nested within a class scope).
  • Definition edges $E_{def}\subseteq S\times D$, representing that a definition $d$ is contained within a scope $s$.
  • Reference edges $E_{ref}\subseteq S\times R$, representing that a reference $r$ occurs within a scope $s$.
  • Binding edges $E_{bind}\subseteq R\times D$, representing that a reference $r$ resolves to a definition $d$ according to the language’s scoping rules. These edges are established during the resolution process.
Thus, $V=S\cup D\cup R$ and $E=E_{scope}\cup E_{def}\cup E_{ref}\cup E_{bind}$.
B.2 Dependency Resolution Process
For each function candidate $F$, we identify the set of all identifier references $R_{F}\subseteq R$ occurring within its body scope $s_{F}$. For each reference $r\in R_{F}$, the dependency resolution process attempts to find a corresponding definition $d\in D$ such that a binding edge $(r,d)\in E_{bind}$ can be established by traversing the Scope Graph $G$ outwards from $s_{F}$ according to the language’s lexical scoping rules.
A reference $r\in R_{F}$ is classified as an unresolved reference if no definition $d\in D$ is found within the analysis scope of $G$ that $r$ can legally bind to. Let $U_{F}\subseteq R_{F}$ be the set of all unresolved references for function $F$. These unresolved references $U_{F}$ represent the external dependencies of function $F$.
The Scope Graph approach offers conceptual language-agnosticism as the graph structure and resolution mechanism are based on universal programming concepts (scopes, bindings), abstracted from specific syntax by the AST input from Tree-sitter. Language-specific scoping rules are encoded in how the resolution traversal and binding edges $E_{bind}$ are determined.
B.3 SC/WSC Classification based on Dependencies
Based on the set of unresolved references $U_{F}$ and the function’s import statements, we classify function $F$:
Let $L_{allowed}$ be the predefined set of allowed common external libraries. For each unresolved reference $r\in U_{F}$, we attempt to determine its origin based on the function’s import statements and knowledge of library APIs. A reference $r$ is considered resolved to an allowed library $l\in L_{allowed}$ if its name $id(r)$ corresponds to an identifier provided by library $l$, and library $l$ is correctly imported or accessible within the scope of function $F$.
Let $U_{F,allowed}\subseteq U_{F}$ be the subset of unresolved references that resolve to identifiers within libraries in $L_{allowed}$.
  • Self-Contained (SC): Function $F$ is classified as SC if and only if its set of unresolved references is empty, i.e., $U_{F}=\emptyset$. This means all identifiers are defined locally or are language built-ins.
  • Weakly Self-Contained (WSC): Function $F$ is classified as WSC if $U_{F}\neq\emptyset$ and all unresolved references resolve to allowed libraries, i.e., $U_{F}=U_{F,allowed}$.
  • Discarded: Function $F$ is discarded if $U_{F}\neq\emptyset$ and $U_{F,allowed}\subsetneq U_{F}$. This means there are unresolved references that are not from allowed libraries.
This systematic, formalized approach, leveraging the Scope Graph representation, provides a precise and robust method for controlling dependencies and classifying functions, which is essential for the reliability and scalability of the Code2Bench.
Appendix C Program Analysis for Testability and Complexity
This appendix provides further details on the program analysis techniques employed in the Function Filtering stage (Section 2) to ensure candidate functions are functionally testable and represent non-trivial coding challenges. Our analysis focuses on properties derived from the Control Flow Graph (CFG) and the structural complexity of the function implementation.
C.1 Control Flow Analysis for Testability
To identify functions amenable to automated testing and output verification, we perform Control Flow Graph (CFG) analysis. For a given candidate function $f$, its CFG represents all possible execution paths through the function’s code. Nodes in the CFG correspond to basic blocks of code (sequences of instructions executed sequentially), and directed edges represent potential control transfers between these blocks (e.g., branches, loops, function calls).
We analyze the structure of the CFG to filter out functions that inherently lack verifiable output. Specifically, we identify functions where:
  • The CFG contains no paths leading to a return statement. Such functions typically perform actions (e.g., printing to console, modifying global state) without providing a value that can be easily captured and compared against an expected output in an automated differential testing setup.
  • All paths leading to a return statement return only constant values, or values derived solely from constants without any dependency on input parameters or complex intermediate computations. While these functions have a return value, their behavior is trivial and does not require probing with diverse inputs.
Functions matching these criteria are excluded from the candidate pool, as they are either difficult to test functionally in isolation or do not represent meaningful code generation tasks for evaluating LLM capabilities beyond simple retrieval.
C.2 Complexity Assessment via Cyclomatic Complexity
To focus the benchmark on tasks that require models to generate non-trivial code logic, we assess the structural complexity of each candidate function using Cyclomatic Complexity ($CC$) [31]. Cyclomatic Complexity is a quantitative measure of the number of linearly independent paths through a program’s source code. It is calculated based on the CFG of the function using the formula:
CC=EN+2PCC=E-N+2P
where:
  • $E$ is the number of edges in the CFG.
  • NN is the number of nodes in the CFG.
  • $P$ is the number of connected components in the CFG (for a single function, $P=1$).
Therefore, for a single function, CC=EN+2CC=E-N+2. CCCC is strongly correlated with the number of decision points (e.g., if, while, for, case) in the code, providing an estimation of the code’s logical complexity.
We employ $CC$ as a filter to select functions that fall within a desirable complexity range, avoiding tasks that are either too simple to be challenging or excessively complex to be solvable or reliably testable as isolated benchmark instances. Specifically, we define a range $CC_{min},CC_{max}]$ (e.g., $[2,10]$) and include only functions whose calculated $CC$ falls within this range.
  • Functions with $CC<CC_{min}$ (e.g., $CC<2$) typically represent very simple linear code or trivial control flow structures that offer little challenge.
  • Functions with $CC>CC_{max}$ (e.g., $CC>10$) may indicate highly complex logic, potentially involving deep nesting, numerous branches, or intricate loops, which can be challenging for LLMs to generate correctly and for automated tests to cover exhaustively. Furthermore, such high complexity in a real-world function might often be coupled with complex, uncontrolled dependencies.
Appendix D Semantic Filtering and Difficulty Assessment
This appendix provides additional technical details regarding the LLM-based Semantic Filtering process used in the Candidate Filtering stage (Section 2.1) to refine the candidate function set through semantic assessment. While the main text outlines the motivation and overall structure, here we present in detail the prompts employed during this stage of semantic filtering.
Table 6: Prompt Template for Self-Contained Ground-Truth Filter in Code2Bench
# System
## Task Description
You are an expert in the field of coding, tasked with determining whether a given Python function is suitable for generating an instruction (question).
The function will be analyzed based on its characteristics, functionality, and adherence to specific criteria. If the function meets the criteria, it is deemed suitable; otherwise, it is not.
## Criteria for Suitability To determine whether a function is suitable for generating an instruction, consider the following criteria:
### 1. Function Parameters
- **Basic Types Only**: The function’s parameters must be basic types (e.g., ‘int‘, ‘float‘, ‘str‘, ‘list‘, ‘dict‘, etc.).
- If the parameters include custom classes or unknown types, the function is **not suitable**.
- If the function relies on methods or attributes of unknown objects, the function is **not suitable**.
### 2. Function Complexity
- **Meaningful Complexity**: The function should provide a meaningful test of the model’s capabilities.
- If the function is very long but its logic is trivial, it is **not suitable**.
- If the function is too simple (e.g., basic business logic), it is **not suitable**.
- If the function lacks clear logic or purpose, it is **not suitable**.
### 3. Side Effects and Dependencies
- **No Side Effects**: The function should not have side effects (e.g., modifying global variables, writing to files, etc.).
- **No External Imports**: The function should not import other modules or depend on external libraries.
## Output Format If the function is **suitable**, return:
“‘json
"Suitable": true,
"Reason": "The function meets all criteria for generating an instruction."
If the function is not suitable , return:
“‘json
"Suitable": false,
"Reason": "The reason why the function is not suitable for generating an instruction."
## Examples
## Note
- Ensure that your analysis is thorough and considers all aspects of the function.
- Provide clear and concise reasoning for your decision. - Only return the Json.
# User
Please check the last result:
[Last Result]
Error response:
[Error Response]
[Function Message]
Table 7: Prompt Template for Weakly Self-Contained Ground-Truth Filter in Code2Bench
# System
You are an expert in Python coding, tasked with determining whether a given Python function meets the requirements below for generating a benchmark. The function is weakly self-contained, meaning it depends only on Python standard libraries or specific external libraries (e.g., numpy, re, pandas) and no custom modules. You will analyze the function based on its characteristics, functionality, and adherence to specific criteria. If the function meets the criteria, it is deemed suitable; otherwise, it is not.
## Criteria for Suitability
To determine whether a function is suitable, consider the following criteria:
### 1. Function Parameters
- **Basic and Library Types**: Parameters must be basic Python types (e.g., ‘int‘, ‘float‘, ‘str‘, ‘list‘, ‘dict‘) or types from standard/external libraries (e.g., ‘numpy.ndarray‘, ‘re.Pattern‘).
- If the parameters’ type is missing, but you can infer it from the code, it is **suitable**.
- If the function relies on methods or attributes of unknown objects, it is **not suitable**. But if the function uses lib types (e.g., ‘numpy.ndarray‘, ‘pandas.DataFrame‘), it is **suitable**.
### 2. Function Complexity
- **Meaningful Complexity**: The function should provide a meaningful test of the model’s capabilities, with clear logic and purpose.
- If the function is overly long but trivial (e.g., repetitive assignments), it is **not suitable**.
- If the function is too simple (e.g., basic getter/setter), it is **not suitable**.
### 3. Domain Knowledge
- **General Applicability**: The function should not require highly specialized domain knowledge to understand or implement.
- If the function is tied to a specific, non-generalizable application (e.g., proprietary system logic), it is **not suitable**.
### 4. Property-Based Testing
- **Constructible Inputs**: The function should allow generating random inputs for property-based testing to verify its behavior.
## Output Format
If the function is **suitable**, return:
“‘json
"Suitable": true,
"Reason": "The reason why the function meets all criteria for generating a benchmark."
If the function is not suitable, return:
“‘json
"Suitable": false,
"Reason": "The reason why the function is not suitable for generating a benchmark."
## Examples
## Note
- Provide clear and concise reasoning for the decision.
- If the function meets our standards but is missing imports, it is still suitable.
- If the function only uses ‘typing‘ for type hints, it is not suitable.
- Only return the JSON output in the specified format.
# User
Please check the last result:
[Last Result]
Error response:
[Error Response]
[Function Message]
Table 8: Prompt Template for Weakly Self-Contained Difficulty Assessment in Code2Bench
# System
## Task Description
You are an expert code analyst tasked with assessing the difficulty level of a weakly self-contained Python function. A weakly self-contained function depends only on Python standard libraries or specific external libraries (e.g., NumPy, pandas, re) and no custom modules. Assume the function is valid and suitable for analysis. Assign a difficulty level of "Easy", "Medium", or "Hard" based on the complexity of its logic, structure, required concepts, and cognitive load to understand.
## Criteria for Difficulty Assessment ### Easy Difficulty - **Logic**: Very simple, minimal or no branching, single loop or direct parameter use.
- **Structure**: Short, linear, immediately clear control flow.
- **Concepts**: Basic Python constructs (variables, operators, lists, strings) or simple library calls (e.g., Counter from collections, basic re matching, pandas filtering).
- **Cognitive Load**: Minimal; purpose and execution are obvious at a glance.
- **Example**:
“‘python
from collections import Counter
def count_word_frequencies(text: str) -> dict[str, int]:
"""Count the frequency of each word in a text string.
Args:
text: Input string containing words.
Returns:
Dictionary mapping words to their frequency.
words = text.lower().split()
return dict(Counter(words))
### Medium Difficulty - Logic: Moderate complexity, with loops, conditions, or data transformations (e.g., - filtering, sorting, deduplication).
- Structure: Traceable control flow, possibly nested loops or multiple steps, moderate length.
- Example:
Hard Difficulty
- Logic: Complex, with nested loops, intricate transformations, non-trivial algorithms (e.g., multi-dim aggregation, complex grouping), or subtle edge cases.
- Structure: Dense or multi-step control flow, significant state management.
Example:
## Output Format Return ONLY a JSON object containing the assessed difficulty level:
“‘json
"Difficulty": "Easy/Medium/Hard"
## Note
- weakly self-contained function depends only on Python standard libraries or specific external libraries (e.g., NumPy, pandas, re) and no custom modules.
- Focus on logic and structure, not the library’s complexity (e.g., simple Counter usage is Easy, complex NumPy array ops are Hard).
- Analyze the function’s code, docstring, and logic thoroughly.
- Only return the JSON output in the specified format.
# User
[Function Message]
Table 9: Prompt Template for Self-Contained Difficulty Assessment in Code2Bench
# System
## Task Description
You are an expert code analyst. Your task is to assess the difficulty level of the provided Python function. Assume the function is generally valid and suitable for analysis. Assign a difficulty level of "Easy", "Medium", or "Hard" based on the complexity of its logic, structure, required concepts, and cognitive load to understand.
## Criteria for Difficulty Assessment
### Easy Difficulty
- Logic: Straightforward, minimal branching (simple if/else), possibly a single simple loop. Direct use of parameters.
- Structure: Typically short, linear control flow. Easy to follow step-by-step.
- Concepts: Relies on fundamental programming constructs (variables, basic operators, standard data types, simple function calls).
- Cognitive Load: Low; the function’s purpose and execution are immediately apparent.
- Example:
“‘python
def parse_message(message: str) -> str:
if message is None:
return ""
message = message.strip().lower()
# Simple string checks and manipulations
if not message.startswith(("run-slow", "run_slow", "run slow")):
return ""
message = message[len("run slow") :]
while message.strip().startswith(":"):
message = message.strip()[1:]
return message
### Hard Difficulty
- Logic: Moderate complexity. May involve nested loops, multiple non-trivial conditions, manipulation of data structures (e.g., iterating through lists/dicts with transformations), implementing a common simple algorithm, or tracking state across iterations.
- Structure: Control flow is more involved but still reasonably traceable. Function length might be moderate.
- Example:
### Hard Difficulty
- Logic: Complex logic. Might involve recursion, implementing non-trivial algorithms.
- Structure: Can have nested structures, complex control flow, significant state management, or rely on clever interactions between code parts. May not be long but could be dense.
- Example:
## Output Format
Return ONLY a JSON object containing the assessed difficulty level:
"Difficulty": "Easy/Medium/Hard"
# User
[Function Message]
Appendix E Property-Based Testing
This appendix provides additional technical details regarding the Property-Based Testing (PBT) process used in the Benchmark Instance Construction stage (Section 2.2 ) to generate rigorous test cases for Code2Bench. While it describes the core principles, here we elaborate on the technical implementation.
Our PBT approach is centered around defining strategies for generating diverse, valid inputs for a given function and verifying a core property against the ground truth implementation.
E.1 Strategy Building and Input Synthesis
The Strategy Builder component analyzes the function’s signature, parameter types, and inferred constraints from type hints, docstrings, and static analysis of the ground truth code. It then leverages a PBT library (e.g., Hypothesis for Python) to compose generation strategies for each function parameter. These strategies are designed to explore a wide range of valid inputs, including typical values, edge cases (e.g., empty lists, zero, maximum/minimum values), and combinations of different input types within complex structures (e.g., lists of dictionaries, tuples of specific types). The process is constraint-aware, ensuring generated inputs adhere to inferred conditions.
For example, for a function taking a list of integers ‘def process(data: list[int])‘, the strategy might generate lists of varying lengths, containing both positive and negative integers, zeros, and potentially boundary values like ‘sys.maxint‘. For a function taking a string with specific format requirements, the strategy would be built to generate strings adhering to that format.
E.2 Property Definition and Verification
The core property we verify for generated code is functional equivalence with the ground truth implementation. For every input $x_{i}$ generated by the strategies, we compute the expected output $y_{i}=f_{GT}(x_{i})$, where $f_{GT}$ is the ground truth function. The PBT framework then requires that for any generated input $x_{i}$, the output of the generated code $f_{LLM}(x_{i})$ must equal $y_{i}$. Any input $x_{i}$ where $f_{LLM}(x_{i})\neq y_{i}$ constitutes a test case failure, and the PBT framework can then attempt to "shrink" $x_{i}$ to find a minimal failing input.
E.3 Ensuring Test Rigor and Coverage
To ensure the generated test suites are truly rigorous, we incorporate a quality control step based on test coverage. After generating a suite of $(x_{i},y_{i})$ pairs using PBT strategies, we execute these test cases against the ground truth implementation itself. We use code coverage tools (e.g., “coverage.py” for Python) to measure the branch coverage achieved by the generated test suite on the ground truth code. Only test suites that achieve a high coverage threshold (e.g., 100% average branch coverage) are accepted and included in the final benchmark instance. This filtering step is crucial: even if a strategy can generate many inputs, if those inputs don’t exercise the complex branching logic of the function, the resulting test suite is not rigorous enough to effectively verify model implementations.
E.4 Code Example (Python/Hypothesis)
Below is a simplified illustrative example using the Python Hypothesis library to demonstrate how strategies and properties are defined.
import hypothesis.strategies as st
from hypothesis import given, settings, Verbosity
[…]
# Strategy for generating arbitrary strings
string_strategy = st.text()
[…]
return s[::-1]
[…]
# Hypothetical LLM output, might have bugs
return "".join(reversed(s))
[…]
@given(s=string_strategy) # Use the defined strategy to generate input s
@settings(max_examples=1000) # Example settings for testing
def test_reverse_property(s):
# Property: Reversing twice returns the original string
assert $reverse_string($s)) == s # This property tests the function against itself
[…]
expected_output = ground_truth_reverse(s)
actual_output = llm_generated_reverse(s)
assert actual_output == expected_output # This property tests LLM output against GT
[…]
This example illustrates the basic concept of defining strategies for input generation and asserting properties about the function’s behavior for these inputs. In the Code2Bench pipeline, these principles are automated and scaled to generate comprehensive test suites for a wide variety of functions extracted from real-world code.
E.5 Ensuring Test Case Validity and Cross-language Portability
This section elaborates on how Code2Bench ensures the generated test cases are both valid for evaluation and portable across different target languages for SC tasks, crucial aspects of our rigorous testing methodology.
Ensuring Validity via Ground Truth Validation
A key step in test case generation is ensuring that the inputs generated by PBT strategies are valid and do not cause unexpected errors in the ground truth implementation. While strategies aim to generate diverse inputs, including edge cases, some inputs might be fundamentally invalid according to the function’s inherent constraints or assumptions (e.g., negative values for a parameter that must be non-negative, or strings in an unexpected format). We execute the ground truth function with each generated input. If the ground truth throws an expected exception (e.g., “ValueError” for invalid input according to the function’s contract) or any unexpected runtime error, that input is discarded. Only inputs that the ground truth can process without crashing, producing a valid output, are included in the final test suite. This process, illustrated in the code example below, guarantees that all test cases in the benchmark are meaningful and that failures of the LLM’s generated code are due to functional incorrectness rather than invalid inputs that crash the ground truth.
Ensuring Cross-language Portability for SC Tasks
For Self-Contained (SC) tasks, the generated test suites are designed to be reusable for evaluating LLMs across multiple programming languages (Python, Java, Go, TS, JS). To achieve this, the Strategy Builder component enforces constraints on the generated inputs to ensure they use basic data types (integers, floats, strings, lists, dictionaries, booleans) with ranges safely representable across these common target languages. For instance, instead of allowing Hypothesis to generate Python’s arbitrary precision integers, explicit “min_value” and “max_value” bounds are set for integer strategies to fit within standard 32-bit or 64-bit integer limits common in Java, Go, and TypeScript. Similarly, constraints are applied to float ranges and the complexity of nested structures. This focus on portable basic types and bounded ranges ensures that the test case data can be correctly loaded and processed by test runners in different languages without issues like integer overflow or incompatible data structures, making the evaluation truly cross-lingual for SC tasks.
Code Example
Below is a simplified code example illustrating the core PBT generation logic, including calling the ground truth and handling exceptions to ensure test case validity.
import hypothesis.strategies as st
from hypothesis import settings, given
# import other necessary libraries
[…]
# Might raise exceptions for invalid inputs
def example_ground_truth(input1, input2):
# implementation logic
pass
[…]
def strategy_for_input1():
# strategy definition for input1
pass
[…]
# strategy definition for input2
# Ensure ranges are portable for SC tasks!
pass
[…]
# @settings(…)
@given(
input1=strategy_for_input1(),
input2=strategy_for_input2()
def test_generate_valid_cases(input1, input2):
# 1. Validate input and Compute expected output using the ground truth
try:
# Call ground truth with generated inputs
expected_output = example_ground_truth(input1, input2)
except Exception as e: # Catch any exception from ground truth
# Discard inputs that caused errors in the ground truth
return # Hypothesis discard mechanism
[…]
# This pair is collected and saved for the benchmark.
test_case = {"Inputs": {"input1": input1, "input2": input2}, "Expected": expected_output}
collection_list.append(test_case)
pass # Placeholder for collection logic
[…]
This example code snippet demonstrates the core mechanism of validating generated inputs by running the ground truth and collecting only those inputs that result in a valid output, thereby ensuring test case validity.
Appendix F Testcase Runner Generation
This appendix provides a detailed presentation of the prompts used for generating test runners (Section 2). Since test runners are inherently language-specific, we designed tailored prompts for each programming language. The corresponding prompts are listed in Tables 10 through 13.
Table 10: Prompt Template for TS Testcase Runner Generation in Code2Bench
# System
## Task Description
As an expert TypeScript developer specializing in test case generation and function signature translation, your task is to generate a function signature based on a Python function and its test cases.
### Requirements
1. **Function Signature**:
- Match the Python function’s input and output types.
- Use export function declarations, include a // TODO: Implement comment inside the function.
- Output in ‘<signature>‘ tags.
2. **Special Considerations**:
- Use explicit and specific TypeScript types, avoiding any or object for type safety.
- Handle complex inputs (e.g., arrays, tuples, nested structures) precisely.
- Return only ‘<signature>‘ sections, without additional explanations.
## Input Format
- **Python Function**: A Python function to be tested, including its signature and implementation, provided as code.
## Output Format
“‘plaintext
<signature>
[TypeScript function signature]
</signature>
## Examples
### Example 1: fast_format_html
#### Input
Python Function:
“‘python
def fast_format_html(html_string):
indent = 0
indent_str = " "
formatted = []
[…]
for part in parts:
if not part.strip():
continue
if part.startswith("</"):
indent -= 1
formatted.append(indent_str * indent + part)
elif part.startswith("<") and part.endswith("/>"):
formatted.append(indent_str * indent + part)
# User
The previously generated runner code:
[Runner Code]
The previously generated runner code resulted in the following error during execution:
[Error Message]
[Function Message]
Table 11: Prompt Template for Java Testcase Runner Generation in Code2Bench
# System
## Task Description
As an expert Java developer specializing in test case generation and function signature translation, your task is to generate a Java test file and function signature based on a Python function and its test cases. The test file will load test cases from a JSON file and execute tests using a provided ‘Helper.deepCompare‘ method.
### Requirements
1. **Java Test File**:
- Generate a complete, executable Java test file in the ‘p0‘ package.
2. **Function Signature**:
- Provide only the function signature in the ‘p0.Tested‘ class.
3. **Special Considerations**:
- **Type Safety**:
- Ensure ‘TestCase‘ fields exactly match JSON keys and types (e.g., ‘lines‘ → ‘List<String>‘, ‘line_index‘ → ‘int‘).
4. **Type Definition Rules**
- Follow these rules to determine where to define types:
| Usage Scenario | Location | Example
| Used in function signature | ‘tested.java‘ | ‘public static class TagInfo {}‘ |
| Used in both | ‘tested.java‘ | Shared types always in implementation | ## Input Format
- **Test Cases JSON**: A JSON array of test cases, provided as ‘{testcases_str}‘.
- **Python Function**: A Python function to be tested, including its signature and implementation, provided as code. ## Output Format
“‘plaintext
<code>
[Java test file]
</code>
<signature>
[Java function signature]
</signature> ## Examples
## Note
- Ensure the generated Java test file is complete and executable.
- The function signature should be a valid Java function signature that matches the Python function’s behavior.
- Only return the Java test file in ‘<code>‘ and the function signature in ‘<signature>‘ tags.
# User
The previously generated runner code:
[Runner Code]
The previously generated runner code resulted in the following error during execution:
[Error Message]
[Function Message]
Table 12: Prompt Template for Go Testcase Runner Generation in Code2Bench
# System
## Task Description
As an expert Go developer, generate two files from Python function definitions and JSON test cases:
1. ‘runner_test.go‘ - Test execution logic only
2. ‘tested.go‘ - Implementation signature with all required types
### Strict Requirements
1. **File Separation Rules**:
2. **Error Handling Rules**:
- If Python function can raise exceptions:
→ Go function MUST return ‘(resultType, error)‘
→ Test MUST verify no error occurs (since test cases are all valid inputs)
- If Python has no error cases:
→ Go function returns only ‘resultType‘
3. **Type Definition Rules**:
| Usage Scenario | Location | Example |
| Used in function signature | ‘tested.go‘ | ‘type TagInfo struct{}‘ |
| Only in test JSON structure | Inline in TestCase | ‘Inputs struct{ Temp float64 }‘ |
4. **Test Validation Logic**:
- For functions returning ‘(result, error)‘:
### Output Format
“‘plaintext
<code>
// runner_test.go
[Complete test file content]
</code>
<signature>
// tested.go
[Complete implementation file content]
</signature>
## Examples
## Note
- Ensure the generated runner_test.go file is complete and executable.
- The function signature should be a valid Go function signature that matches the Python function’s behavior.
- Only return the runner_test.go file in ‘<code>‘ and the tested.go in ‘<signature>‘ tags.
# User
The previously generated runner code:
[Runner Code]
The previously generated runner code resulted in the following error during execution:
[Error Message]
[Function Message]
Table 13: Prompt Template for Weakly Self Contained Testcase Runner Generation in Code2Bench
# System
## Task Description
You are an expert Python developer specializing in property-based testing and test case execution. Your task is to generate a **complete and executable Python script** that loads test cases from a JSON file generated by a Hypothesis-based Testcase Generator and re-runs the test logic to verify the behavior of a function under test (‘func1‘) against a ground truth function (‘func0‘). The functions depend only on standard libraries or specific external libraries (e.g., NumPy, re) and no other custom modules. The script will:
1. Load test cases from the JSON file (‘test_cases.json‘) containing 500 test cases, each with a ‘"Inputs"‘ dictionary mapping ‘func0‘’s argument names to JSON-serializable values.
2. Re-run the test logic by calling ‘func0‘ (ground truth) and ‘func1‘ (under test) with the loaded inputs, comparing their outputs via differential testing.
3. Compare outputs using:
- For basic types and their combinations (‘int‘, ‘float‘, ‘str‘, ‘list‘, ‘dict‘, etc.), use ‘deep_compare‘ from the ‘helper‘ module.
- For third-party library types (e.g., ‘numpy.ndarray‘, ‘tuple‘ of ‘numpy.ndarray‘), use library-provided comparison functions (e.g., ‘np.allclose‘) or custom logic if none is provided.
4. Report test results, indicating whether each test case passes or fails, with detailed failure information (inputs, expected output, actual output).
## Input The input is a Python Testcase Generator script that includes:
1. The ground truth function ‘func0‘, its dependencies (e.g., ‘numpy‘, ‘re‘), and implementation.
2. Hypothesis strategies and ‘@example‘ decorators defining input generation logic.
3. A test function (‘test_<function_name>‘) that generates and saves 500 test cases to ‘test_cases.json‘, each containing only ‘"Inputs"‘.
Provided in ‘<Testcase Generator>‘ tags.
## Output Generate a complete, executable Python script that:
## Example
## Note - Focus on Loading and Re-testing: Load test cases from test_cases.json and verify func1 against func0 using differential testing.
- Preserve Input Format: Ensure inputs match func0’s signature, converting JSON-serialized inputs (e.g., list to np.ndarray) as needed.
- Output Comparison:
- Use deep_compare from helper for basic types and combinations (int, float, str, list, dict, etc.).
- Use library-provided comparisons (e.g., np.allclose for numpy.ndarray) for third-party library types, or custom logic if none is provided.
- Executable Code: The script must be complete, self-contained, and executable.
- Differential Testing: Since test cases contain only inputs, compute expected outputs by calling func0 and compare with func1’s outputs.
- External Libraries: Include imports for func0’s dependencies (e.g., numpy, re) and handle their data types (e.g., np.ndarray).
# User
The previously generated runner code:
[Runner Code]
The previously generated runner code resulted in the following error during execution:
[Error Message]
[Function Message]
Appendix G Instruction Generation
Table 14: Prompt Template for Weakly Self Contained Instruction Generation in Code2Bench
# System
You are a python programming expert who is refining docstrings in existing programs. You will be given a python function in a python file with an existing (possibly underspecified) docstring with corresponding some input-output examples extracted. Your goal is to refine the associated docstring by making it more informative, precise and complete without adding verbosity or detailed programming logic to the docstring. When there is a docstring, the docstring is used to evaluate the code generation capabilities of a model.
The docstring should particularly describe the format and types of the expected inputs and output as well as the behavior of the function. Do not guess outputs for functions. Finally, do not throw away existing details from the docstrings and only insert content you are sure about. Do NOT have repeated content in the docstring and ONLY describe the high-level function behavior without going into implementation details.
## Requirements
1. **Core Description Fidelity**: The docstring must accurately reflect the function’s behavior and describes the task this function solves. **Pay close attention to the sequence of checks, conditions, and resulting actions within the code.
2. ** Highlight any **special rules** that affect the model’s correct understanding of the function’s behavior, such as:
- Recursive behavior
- Merging, flattening, filtering, transformation logic
- Edge cases or type-specific handling
- Magic numbers or constants
3. **Docstring Refinement**: If the function already has a docstring, integrate and refine its content to meet these requirements. Do not discard existing accurate information.
4. **Conditional Example Handling**:
You should judge whether to include examples based on the original docstring’s content:
- **If the original docstring contains an ‘Examples‘ section**:
- Preserve all original examples **verbatim** in the final docstring’s ‘Examples‘ section.
- Format them clearly in Language-Agnostic(e.g., showing input and expected output).
- Do **not** add or modify examples from the ‘Example Usages‘ data.
## Input “‘python
{ground_truth_function_code}
### Output Format - Return the docstring in <docstring> tags following the Google-style format.
- Include the function signature in <signature> tags, with a TODO placeholder for the implementation.
## Note
- Only add examples in Docstring when the function already has an ‘Examples‘ section in the docstring. Do not add examples from the ‘Example Usages‘ data if the original docstring does not contain ‘Examples‘.
- If the signature has type hints, import nessesary types from the standard library (e.g., ‘from typing import List, Dict‘) in signature.
# User
<function>
[Function Code]
</function>
<example usages>
[Example Usage]
</example usages>
Table 15: Prompt Template for Self Contained Instruction Generation in Code2Bench
# System
You are a programming documentation architect specializing in creating precise, implementation-agnostic specifications. Generate a docstring that enables accurate reimplementation in any programming language. When there is a docstring, the docstring is used to evaluate the code generation capabilities of a model.
## Requirements
1. **Core Description Fidelity**: The docstring must accurately reflect the function’s behavior and describes the task this function solves. **Pay close attention to the sequence of checks, conditions, and resulting actions within the code.
2. Highlight any **special rules** that affect the model’s correct understanding of the function’s behavior, such as:
- Recursive behavior
- Edge cases or type-specific handling
- Magic numbers or constants
- Special settings that may affect the difficulty for others to correctly implement functions based on docstrings. e.g., the Ground Truth function may add some special string at the end of the result, so the docstring should mention this case, otherwise, the model may not be able to implement the function correctly.
2. **Language-Agnostic Terminology**: Use universal concepts for types and logic.
- Describe parameters and return values using **conceptual types** (e.g., "an integer", "a boolean value", "a sequence of numbers", "a text string") instead of language-specific type hints (‘int‘, ‘bool‘, ‘list‘, ‘str‘).
- Describe operations conceptually (e.g., "checks if X contains Y", "iterates over the elements", "applies a function to each element") rather than Python built-ins (‘s1.find("’")‘, ‘s1.replace‘).
3. **Docstring Refinement**: If the function already has a docstring, integrate and refine its content to meet these requirements. Do not discard existing accurate information.
4. **Conditional Example Handling**:
- **If the original docstring contains an ‘Examples‘ section**:
- Preserve all original examples **verbatim** in the final docstring’s ‘Examples‘ section.
- Format them clearly in Language-Agnostic(e.g., showing input and expected output).
- Do **not** add or modify examples from the ‘Example Usages‘ data.
## **Input Structure**
“‘python
{ground_truth_function_code}
### Example Usages:
{example_usage_data}
(Note: Example Usages will be provided in a format like input/output pairs.)
## **Output:**
Only return the docstring content in <docstring> tags and the function signature in the <signature> tags. The docstring should be enclosed in triple double quotes (‘"""Docstring goes here"""‘). The function signature should be formatted in “‘python‘ code block with the function name and a TODO comment indicating where the implementation should go.
## Note:
- Only add examples in Docstring when the function already has an ‘Examples‘ section in the docstring. Do not add examples from the ‘Example Usages‘ data if the original docstring does not contain ‘Examples‘.
- If the signature has type hints, import nessesary types from the standard library (e.g., ‘from typing import List, Dict‘) in signature.
# User
<function>
[Function Code]
</function>
<example usages>
[Example Usage]
</example usages>
Appendix H Benchmark
H.1 Benchmark Details
This subsection provides additional details regarding the construction and properties of the Code2Bench-2505 instance.
Source Code Acquisition and Filtering
The real-world code used for constructing Code2Bench-2505 was sampled from GitHub repositories with at least 500 stars. This criterion was applied to focus on relatively mature and well-maintained projects, increasing the likelihood of finding high-quality and realistic code functions. A comprehensive list of the specific GitHub repositories sampled for Code2Bench-2505 is provided alongside the benchmark data and code release at the public URL (as mentioned in abstract).
LLM Usage in Benchmark Construction
Large Language Models play a role in several stages of the Code2Bench pipeline for constructing benchmark instances. For the construction of Code2Bench-2505, GPT-4o was primarily employed for these tasks within the benchmark construction pipeline. It is crucial to emphasize that GPT-4o was strictly excluded from the set of models evaluated in the Evaluation section.
Quality Assurance and Release Process
Prior to the official release of any Code2Bench benchmark instance, a rigorous quality assurance process is performed. This includes a dry run using a designated, capable LLM. Given that the testcase runners for Python, JavaScript (JS), and TypeScript (TS) are primarily constructed using template-based methods, their inherent correctness is highly reliable. Consequently, the dry run specifically emphasizes validating the testcase runners for Java and Go, as well as the WSC Python runner
H.2 Benchmark Task Examples
This appendix provides examples of representative benchmark instances from Code2Bench-2505. Each example showcases a complete task, including the instruction provided to the Large Language Model (LLM), the ground truth implementation from which the task was derived, the Property-Based Testing (PBT) script used for generating comprehensive test cases, and the test runner script for evaluating the LLM’s generated code.
H.2.1 SC Python Example: merge_json_recursive
This example demonstrates a Self-Contained (SC) task in Python, requiring the recursive merging of JSON-like objects without external dependencies beyond standard library features.
Task Instruction
def merge_json_recursive(base, update):
"""Recursively merge two JSON-like objects.
[…]
- If both inputs are dictionaries, recursively merge them.
- If both inputs are lists, concatenate them.
- For all other cases, the update value overwrites the base value.
- The base object is left unmodified; a new merged object is returned.
[…]
base: Base JSON-like object (dictionary, list, or primitive value).
update: Update JSON-like object to merge into base.
[…]
A new JSON-like object containing merged content from base and update.
[…]
Input: base = {"a": 1}, update = {"a": 2}
Output: {"a": 2}
[…]
Output: [1, 2, 3, 4]
[…]
Output: {"a": {"b": 1, "c": 2}}
# TODO: Implement this function
pass
Testcase Generator
from hypothesis import settings, given, Verbosity, example
from hypothesis import strategies as st
import json
import os
import atexit
import copy
[…]
TEST_CASE_DIR = os.path.abspath("test_cases")
os.makedirs(TEST_CASE_DIR, exist_ok=True)
TEST_CASE_FILE = os.path.join(TEST_CASE_DIR, "test_cases.json")
generated_cases = []
stop_collecting = False # Global flag to control case collection
[…]
def merge_json_recursive(base, update):
if not isinstance(base, dict) or not isinstance(update, dict):
if isinstance(base, list) and isinstance(update, list):
return base + update
return update
[…]
for key, value in update.items():
if key in merged:
merged[key] = merge_json_recursive(merged[key], value)
else:
merged[key] = value
[…]
json_strategy = st.recursive(
st.one_of([
st.integers(),
st.floats(allow_nan=False, allow_infinity=False),
st.text(st.characters(whitelist_categories=(’L’, N’, P’, S’, Z’))),
st.booleans()
lambda children: st.one_of(
st.lists(children, max_size=5),
st.dictionaries(st.text(st.characters(whitelist_categories=(’L’, N’)), max_size=5), children, max_size=5)
max_leaves=5
[…]
@settings(max_examples=10000, verbosity=Verbosity.verbose, print_blob=True)
@given(base=json_strategy, update=json_strategy)
def test_merge_json_recursive(base, update):
global stop_collecting
if stop_collecting:
return
[…]
update_copy = copy.deepcopy(update)
expected = merge_json_recursive(base_copy, update_copy)
[…]
generated_cases.append({
"Inputs": {"base": base, "update": update},
"Expected": expected
if len(generated_cases) >= 500:
stop_collecting = True
[…]
def save_test_cases():
with open(TEST_CASE_FILE, "w") as f:
json.dump(generated_cases, f, indent=2, ensure_ascii=False)
print(f"✅ Saved {len(generated_cases)} test cases to {TEST_CASE_FILE}")
[…]
Testcase Runner
import json
import os
from tested import merge_json_recursive as func1
from helper import deep_compare
[…]
TEST_CASE_DIR = os.path.abspath("test_cases")
TEST_CASE_JSON_PATH = os.path.join(TEST_CASE_DIR, "test_cases.json")
[…]
if not os.path.exists(TEST_CASE_JSON_PATH):
print(f"JSON file not found: {TEST_CASE_JSON_PATH}")
return []
[…]
with open(TEST_CASE_JSON_PATH, "r") as f:
test_cases = json.load(f)
[…]
for i, case in enumerate(test_cases):
inputs = case["Inputs"]
expected_output = case["Expected"]
[…]
actual_output = func1(**inputs) # Copy matrix to avoid in-place modification
[…]
if not deep_compare(actual_output, expected_output, tolerance=1e-6):
print(f"Test case {i + 1} failed:")
print(f" Inputs: {inputs}")
print(f" Expected: {expected_output}")
print(f" Actual: {actual_output}")
else:
print(f"Test case {i + 1} passed.")
[…]
test_cases = load_test_cases_from_json()
run_tests_with_loaded_cases(test_cases)
H.2.2 SC JavaScript Example: merge_json_recursive
This example demonstrates a Self-Contained (SC) task in JavaScript, requiring the recursive merging of JSON-like objects without external dependencies beyond standard library features.
Task Instruction
* Recursively merge two JSON-like objects.
* The function merges nested structures with the following rules:
* - If both inputs are dictionaries, recursively merge them.
* - If both inputs are lists, concatenate them.
* - For all other cases, the update value overwrites the base value.
* - The base object is left unmodified; a new merged object is returned.
* @param {any} base - Base JSON-like object (dictionary, list, or primitive value).
* @param {any} update - Update JSON-like object to merge into base.
* @returns {any} A new JSON-like object containing merged content from base and update.
* Examples:
* Input: base = {"a": 1}, update = {"a": 2}
* Output: {"a": 2}
* Input: base = [1, 2], update = [3, 4]
* Output: [1, 2, 3, 4]
* Input: base = {"a": {"b": 1}}, update = {"a": {"c": 2}}
* Output: {"a": {"b": 1, "c": 2}}
[…]
// TODO: Implement
Testcase Runner
import fs from fs’;
import path from path’;
import { merge_json_recursive as func } from ’./tested’;
[…]
const rawData = fs.readFileSync(filePath, utf-8’);
const testCases = JSON.parse(rawData);
return testCases;
[…]
const testCases = loadTestCases(’./3/test_cases/test_cases.json’);
[…]
test(‘Case {index}‘, () => {
let base_1 = tc.Inputs.base;
let update_1 = tc.Inputs.update;
[…]
return Math.abs(a - b) <= tolerance;
[…]
if (typeof a === "number" && typeof b === "number") {
return isClose(a, b, tolerance);
} else if (Array.isArray(a) && Array.isArray(b)) {
if (a.length !== b.length) return false;
return a.every((ai, index) => deepCompare(ai, b[index], tolerance));
} else if (
a !== null && typeof a === "object" &&
b !== null && typeof b === "object"
const aKeys = Object.keys(a);
const bKeys = Object.keys(b);
if (aKeys.length !== bKeys.length) return false;
return aKeys.every(k => k in b && deepCompare(a[k], b[k], tolerance));
} else {
return a === b;
H.2.3 SC TypeScript Example: merge_json_recursive
This example demonstrates a Self-Contained (SC) task in TypeScript, requiring the recursive merging of JSON-like objects without external dependencies beyond standard library features.
Task Instruction
* Recursively merge two JSON-like objects.
* The function merges nested structures with the following rules:
* - If both inputs are dictionaries, recursively merge them.
* - If both inputs are lists, concatenate them.
* - For all other cases, the update value overwrites the base value.
* - The base object is left unmodified; a new merged object is returned.
* @param base - Base JSON-like object (dictionary, list, or primitive value).
* @param update - Update JSON-like object to merge into base.
* @returns A new JSON-like object containing merged content from base and update.
* Examples:
* Input: base = {"a": 1}, update = {"a": 2}
* Output: {"a": 2}
* Input: base = [1, 2], update = [3, 4]
* Output: [1, 2, 3, 4]
* Input: base = {"a": {"b": 1}}, update = {"a": {"c": 2}}
* Output: {"a": {"b": 1, "c": 2}}
[…]
// TODO: Implement
Testcase Runner
import fs from fs’;
import path from path’;
import { merge_json_recursive as func } from ’./tested’;
[…]
const rawData = fs.readFileSync(filePath, utf-8’);
const testCases = JSON.parse(rawData);
return testCases;
[…]
const testCases = loadTestCases(’./3/test_cases/test_cases.json’);
[…]
test(‘Case {index}‘, () => {
let base_1: any = tc.Inputs.base;
let update_1: any = tc.Inputs.update;
[…]
return Math.abs(a - b) <= tolerance;
[…]
if (typeof a === "number" && typeof b === "number") {
return isClose(a, b, tolerance);
} else if (Array.isArray(a) && Array.isArray(b)) {
if (a.length !== b.length) return false;
return a.every((ai, index) => deepCompare(ai, b[index], tolerance));
} else if (
a !== null && typeof a === "object" &&
b !== null && typeof b === "object"
const aKeys = Object.keys(a);
const bKeys = Object.keys(b);
if (aKeys.length !== bKeys.length) return false;
return aKeys.every(k => k in b && deepCompare(a[k], b[k], tolerance));
} else {
return a === b;
H.2.4 SC Java Example: merge_json_recursive
This example demonstrates a Self-Contained (SC) task in Java, requiring the recursive merging of JSON-like objects without external dependencies beyond standard library features.
Task Instruction
package p3;
import java.util.Map;
import java.util.List;
[…]
* Recursively merge two JSON-like objects.
* The function merges nested structures with the following rules:
* - If both inputs are dictionaries, recursively merge them.
* - If both inputs are lists, concatenate them.
* - For all other cases, the update value overwrites the base value.
* - The base object is left unmodified; a new merged object is returned.
* @param base Base JSON-like object (dictionary, list, or primitive value).
* @param update Update JSON-like object to merge into base.
* @return A new JSON-like object containing merged content from base and update.
* Examples:
* Input: base = {"a": 1}, update = {"a": 2}
* Output: {"a": 2}
* Input: base = [1, 2], update = [3, 4]
* Output: [1, 2, 3, 4]
* Input: base = {"a": {"b": 1}}, update = {"a": {"c": 2}}
* Output: {"a": {"b": 1, "c": 2}}
public class Tested {
public static Object merge_json_recursive(Object base, Object update) {
// TODO: Implement this function
Testcase Runner
package p3;
[…]
import com.fasterxml.jackson.core.type.TypeReference;
import org.junit.jupiter.params.ParameterizedTest;
import org.junit.jupiter.params.provider.MethodSource;
import java.io.File;
import java.io.IOException;
import java.util.List;
import java.util.Map;
import java.util.stream.Stream;
import static org.junit.jupiter.api.Assertions.assertTrue;
[…]
// TestCase represents the structure of our test cases
static class TestCase {
public Inputs Inputs;
public Object Expected; // Use Object to support multiple types
[…]
public Object base; // Use Object to support Map or List
public Object update; // Use Object to support Map or List
[…]
private static List<TestCase> loadTestCases(String filePath) throws IOException {
ObjectMapper mapper = new ObjectMapper();
File file = new File(filePath);
try {
return mapper.readValue(file, new TypeReference<List<TestCase>>() {
} catch (IOException e) {
throw new IOException("Failed to load test cases from " + filePath + ": " + e.getMessage(), e);
[…]
private static Stream<TestCase> testCases() throws IOException {
return loadTestCases("src/test/java/test_cases/test_cases.json").stream();
[…]
@MethodSource("testCases")
void testMergeJsonRecursive(TestCase tc) {
try {
// Call the function under test
Object actual = Tested.merge_json_recursive(tc.Inputs.base, tc.Inputs.update);
[…]
assertTrue(
Helper.deepCompare(tc.Expected, actual, 0),
String.format(
"Test case failed:tc.Inputs.base, tc.Inputs.update, tc.Expected, actual));
} catch (Exception e) {
throw new RuntimeException("Unexpected exception during test execution", e);
H.2.5 SC Go Example: merge_json_recursive
This example demonstrates a Self-Contained (SC) task in Go, requiring the recursive merging of JSON-like objects without external dependencies beyond standard library features.
Task Instruction
package main
[…]
// The function merges nested structures with the following rules:
// - If both inputs are dictionaries, recursively merge them.
// - If both inputs are lists, concatenate them.
// - For all other cases, the update value overwrites the base value.
// - The base object is left unmodified; a new merged object is returned.
// Params:
// base: Base JSON-like object (dictionary, list, or primitive value).
// update: Update JSON-like object to merge into base.
// Returns:
// A new JSON-like object containing merged content from base and update.
// Examples:
// Input: base = {"a": 1}, update = {"a": 2}
// Output: {"a": 2}
// Input: base = [1, 2], update = [3, 4]
// Output: [1, 2, 3, 4]
// Input: base = {"a": {"b": 1}}, update = {"a": {"c": 2}}
// Output: {"a": {"b": 1, "c": 2}}
func merge_json_recursive(base, update interface{}) interface{} {
// TODO: Implement
Testcase Runner
package main
[…]
"encoding/json"
"fmt"
"os"
"testing"
[…]
type TestCase struct {
Inputs struct {
Base interface{} json:"base"‘
Update interface{} json:"update"‘
} json:"Inputs"‘
Expected interface{} json:"Expected"‘
[…]
func loadTestCases(filePath string) ([]TestCase, error) {
// Read the test cases file
file, err := os.ReadFile(filePath)
if err != nil {
return nil, fmt.Errorf("failed to read test cases: }
[…]
var testCases []TestCase
if err := json.Unmarshal(file, &testCases); err != nil {
return nil, fmt.Errorf("failed to parse test cases: }
[…]
func TestMergeJSONRecursive(t *testing.T) {
// Load test cases from the JSON file
testCases, err := loadTestCases("test_cases/test_cases.json")
if err != nil {
t.Fatalf("Failed to load test cases: }
[…]
for i, tc := range testCases {
t.Run(fmt.Sprintf("Case// Call the function under test
actual := merge_json_recursive(tc.Inputs.Base, tc.Inputs.Update)
[…]
if !DeepCompare(actual, tc.Expected, 0) {
t.Errorf(
Test case Input: base=Expected: Actual: i, tc.Inputs.Base, tc.Inputs.Update, tc.Expected, actual,
H.3 WSC Python Example: calculate_ngram_repetition
This example demonstrates a Weakly Self-Contained (WSC) task in Python. It requires interacting with a function from the standard library (“collections.Counter”) to calculate n-gram repetition in text.
Task Instruction
from collections import Counter
[…]
Calculates the proportion of repeated n-grams in a given text.
[…]
text (str): The input text to analyze, consisting of words separated by spaces.
n (int): The size of the n-grams to generate (e.g., 2 for bigrams, 3 for trigrams).
[…]
float: The proportion of n-grams that are repeated in the text. Returns 0 if no n-grams can be generated.
[…]
ValueError: If n is less than or equal to 0.
# TODO: Implement this function
pass
Testcase Generator
from hypothesis import settings, given, Verbosity, example
from hypothesis import strategies as st
import json
import os
import atexit
import copy
from collections import Counter
[…]
TEST_CASE_DIR = os.path.abspath("test_cases")
os.makedirs(TEST_CASE_DIR, exist_ok=True)
TEST_CASE_FILE = os.path.join(TEST_CASE_DIR, "test_cases.json")
generated_cases = []
stop_collecting = False # Global flag to control case collection
[…]
def calculate_ngram_repetition(text, n):
words = text.split()
ngrams = [tuple(words[i : i + n]) for i in range(len(words) - n + 1)]
ngram_counts = Counter(ngrams)
total_ngrams = len(ngrams)
repeated_ngrams = sum(1 for count in ngram_counts.values() if count > 1)
return repeated_ngrams / total_ngrams if total_ngrams > 0 else 0
[…]
def text_strategy():
return st.text(
alphabet=st.characters(whitelist_categories=(’L’, N’, Z’), min_codepoint=32, max_codepoint=126),
min_size=0, max_size=100
[…]
return st.integers(min_value=1, max_value=5)
[…]
@settings(max_examples=10000, verbosity=Verbosity.verbose, print_blob=True)
@given(text=text_strategy(), n=n_strategy())
@example(text="", n=1)
@example(text="a", n=1)
@example(text="a b c", n=2)
@example(text="a a b b c c", n=2)
@example(text="a b c d e f", n=3)
@example(text="a a a a a a", n=3)
def test_calculate_ngram_repetition(text, n):
global stop_collecting
if stop_collecting:
return
[…]
text_copy = copy.deepcopy(text)
n_copy = copy.deepcopy(n)
[…]
try:
expected = calculate_ngram_repetition(text_copy, n_copy)
except Exception:
return # Skip inputs that cause exceptions
[…]
generated_cases.append({
"Inputs": {
"text": text_copy,
"n": n_copy
[…]
if len(generated_cases) >= 500:
stop_collecting = True
[…]
def save_test_cases():
with open(TEST_CASE_FILE, "w") as f:
json.dump(generated_cases, f, indent=2, ensure_ascii=False)
print(f"✅ Saved {len(generated_cases)} test cases to {TEST_CASE_FILE}")
[…]
Testcase Runner
import json
import os
import copy
from collections import Counter
from helper import deep_compare
from tested import calculate_ngram_repetition as func1
[…]
TEST_CASE_DIR = os.path.abspath("test_cases")
TEST_CASE_JSON_PATH = os.path.join(TEST_CASE_DIR, "test_cases.json")
[…]
def calculate_ngram_repetition(text, n):
words = text.split()
ngrams = [tuple(words[i : i + n]) for i in range(len(words) - n + 1)]
ngram_counts = Counter(ngrams)
total_ngrams = len(ngrams)
repeated_ngrams = sum(1 for count in ngram_counts.values() if count > 1)
return repeated_ngrams / total_ngrams if total_ngrams > 0 else 0
[…]
if not os.path.exists(TEST_CASE_JSON_PATH):
print(f"JSON file not found: {TEST_CASE_JSON_PATH}")
return []
with open(TEST_CASE_JSON_PATH, "r") as f:
test_cases = json.load(f)
return test_cases
[…]
# Use deep_compare for basic types (int, float, str, etc.)
return deep_compare(expected, actual, tolerance=1e-6)
[…]
for i, case in enumerate(test_cases):
inputs = copy.deepcopy(case["Inputs"])
text = inputs["text"]
n = inputs["n"]
[…]
expected_output = calculate_ngram_repetition(text, n)
actual_output = func1(text, n)
[…]
if compare_outputs(expected_output, actual_output):
print(f"Test case {i + 1} passed.")
else:
print(f"Test case {i + 1} failed:")
print(f" Inputs: {inputs}")
print(f" Expected: {expected_output}")
print(f" Actual: {actual_output}")
[…]
test_cases = load_test_cases_from_json()
run_tests_with_loaded_cases(test_cases)
H.4 Benchmark Runner Prompt
Table 16: Prompt Template for Python Benchmark Runner in Code2Bench
# System
You are an expert in the field of coding, helping users write Python code.
## Input
The user provides you with an function signature and docstring, you should generate a Python function based on them.
## Output
“‘python The generated Python code. “‘
## Note
- Only output Python code with possible type import statements but without docstring and any additional information.
# User
[Instruction]
Table 17: Prompt Template for Weakly Self-Contained Benchmark Runner in Code2Bench
# System
You are a highly skilled Python programming expert tasked with implementing a function based on its specification, using the allowed libraries.
Implement the Python function described below. Your implementation should strictly adhere to the behavior specified in the docstring and utilize only the explicitly allowed external libraries.
## Output Format
“‘python The generated Python code. “‘
Provide ONLY the Python code for the function implementation with corrsponding libraries imported. Do not include any additional information or explanations.
# User
[Instruction]
Table 18: Prompt Template for Go Benchmark Runner in Code2Bench
# System
You are an expert in the field of coding, helping users write Go code.
## Input
The user provides you with an function signature and docstring, you should generate a Go function based on them.
## Output “‘go The generated Go code. “‘
## Note
Provide only Go code within a “‘go“‘ code block, including a complete function with package name, necessary imports, and any required type definitions. Do not repeat docstring provided to you. Do not add additional explanations, or a main function.
# User
[Instruction]
Table 19: Prompt Template for Java Benchmark Runner in Code2Bench
# System
You are an expert in the field of coding, helping users write Java code.
## Input
The user provides you with an function signature and docstring, you should generate a Java function based on them.
## Output
“‘java
The generated Java code.
## Note
- Provide only Java code within a “‘java“‘ code block. Include a complete public class named Tested with package name and necessary imports. Do not add a main method or repeat the docstring.
# User
[Instruction]
Table 20: Prompt Template for TypeScript Benchmark Runner in Code2Bench
# System
You are an expert in the field of coding, helping users write TypeScript code.
## Input The user provides you with an function signature and docstring, you should generate a TypeScript function based on them.
## Output
“‘ts
The generated TypeScript code.
## Note
- Only output self-contained TypeScript code without docstring and any additional information.
# User
[Instruction]
Table 21: Prompt Template for JavaScript Benchmark Runner in Code2Bench
# System
You are an expert in the field of coding, helping users write JavaScript code.
## Input
The user provides you with an function signature and docstring, you should generate a JavaScript function based on them.
## Output
“‘js
The generated JavaScript code.
## Note
- Only output JavaScript code without docstring and any additional information.
# User
[Instruction]
Appendix I Detailed Evaluation
I.1 Compute Resources
we detail the compute resources utilized for evaluating the Large Language Models on the Code2Bench-2505 benchmark. The evaluation was conducted on an infrastructure consisting of server-grade machines. Open-source models with fewer than 32B parameters (as listed in Table 2) were served using vLLM on a cluster equipped with NVIDIA GPUs. Specifically, these models were evaluated on machines featuring NVIDIA A100 80GB GPUs. The evaluation environment for these models was containerized to maintain isolation and consistency. For larger open-source models (>= 32B parameters) and all closed-source models, evaluation was performed by accessing their respective official APIs. The compute resources for these API-based evaluations are managed by the model providers and are not under our direct control or knowledge. Therefore, we cannot provide specific details on the underlying hardware, memory, or parallelization used by these providers. The Testcase Runner execution for each task (which involves loading test cases, running the generated code and ground truth, and performing differential testing) was primarily CPU-bound and ran on standard server CPUs, such as Intel(R) Xeon(R) Gold 5218 CPU @ 2.30GHz, featuring 64 logical cores. These machines were equipped with 125 GiB of RAM and SSD storage for the benchmark data and test cases. The total compute time required for the comprehensive evaluation of all 16 models across the 1163 tasks in CODE2BENCH-2505 was substantial. While precise timing varies per model and task, we estimate the total GPU-hours consumed for the open-source model inferences to be approximately 200 GPU-hours. The total CPU-hours consumed for Testcase Runner execution across all models (including running the generated code and ground truth against approximately 500 tests per task per model) is estimated to be approximately 200 CPU-hours.
I.2 Detailed Performance and Failure Analysis
This appendix provides a detailed breakdown of the performance and failure type distribution for each evaluated model across all task types (Self-Contained tasks in Python, JS, TS, Java, Go, and Weakly Self-Contained tasks in Python). The results for each model and task type are presented as percentages of the total number of tasks in that category, broken down into successful completions (Success), completions resulting in execution errors (ExecutionError), and completions failing test cases (TestFailure). The sum of these three percentages for each task type and model is approximately 100%. For clarity, the data is split across multiple tables.
Table 22: Detailed Performance and Failure Type Distribution (Part 1). S=Success, E=Exec. Error, F=Test Failure.
Model Name SC Python SC Go SC Java SC TS SC JS WSC Python
S E F S E F S E F S E F S E F S E F
Doubao-1-5-pro-32k 36.27 6.74 56.99 27.98 28.50 43.52 25.91 10.36 63.73 35.23 5.70 59.07 34.72 5.18 60.10 53.02 9.09 37.89
Gemini-2.5-flash 39.90 7.77 52.33 27.46 28.50 44.04 28.50 16.06 55.44 34.20 7.77 58.03 35.23 7.77 56.99 53.02 8.08 38.90
Gemma-3-27b-it 34.20 5.70 60.10 21.24 30.57 48.19 22.80 15.54 61.66 28.50 10.88 60.62 29.53 10.88 59.59 45.96 8.08 45.96
Mistral-small-3.1 31.61 9.84 58.55 12.95 51.30 35.75 22.80 17.10 60.10 23.83 24.87 51.30 30.57 5.70 63.73 44.44 11.62 43.94
Table 23: Detailed Performance and Failure Type Distribution (Part 2). S=Success, E=Exec. Error, F=Test Failure.
Model Name SC Python SC Go SC Java SC TS SC JS WSC Python
S E F S E F S E F S E F S E F S E F
LLAMA-3.2-3B 18.13 19.69 62.18 3.11 77.20 19.69 6.74 33.16 60.10 7.77 34.20 58.03 9.33 37.82 52.85 23.23 23.74 53.03
LLAMA-4-scout 31.09 7.25 61.66 6.22 78.24 15.54 17.10 29.53 53.37 27.98 16.06 55.96 25.39 10.88 63.73 42.93 11.11 45.96
LLAMA-4-maverick 39.38 6.22 54.40 11.40 66.32 22.28 22.80 19.69 57.51 30.60 30.57 38.83 34.21 23.32 42.47 53.54 7.58 38.89
LLAMA-3.3-70B 37.31 7.25 55.44 15.54 59.07 25.39 24.35 18.65 56.99 27.46 15.03 57.51 27.46 9.33 63.21 45.45 8.08 46.46
Table 24: Detailed Performance and Failure Type Distribution (Part 3). S=Success, E=Exec. Error, F=Test Failure.
Model Name SC Python SC Go SC Java SC TS SC JS WSC Python
S E F S E F S E F S E F S E F S E F
Qwen3-235b-a22b 37.82 6.22 55.96 22.28 39.38 38.34 27.46 13.99 58.55 33.16 8.29 58.55 29.53 6.74 63.73 50.00 9.60 40.40
Qwen3-32b 35.23 7.25 57.51 10.36 48.70 40.93 24.87 14.51 60.62 30.05 7.77 62.18 27.98 7.77 64.25 45.45 11.62 42.93
Qwen2.5-Coder-14b 31.61 8.81 59.59 16.58 55.96 27.46 22.80 12.95 64.25 29.53 9.33 61.14 29.53 9.33 61.14 43.43 16.16 40.40
Qwen2.5-Coder-32b 36.27 8.29 55.44 17.10 59.07 23.83 24.35 12.95 62.69 32.12 6.74 61.14 34.20 10.88 54.92 44.44 14.65 40.91
Table 25: Detailed Performance and Failure Type Distribution (Part 4). S=Success, E=Exec. Error, F=Test Failure.
Model Name SC Python SC Go SC Java SC TS SC JS WSC Python
S E F S E F S E F S E F S E F S E F
Qwen3-8b 28.50 17.62 53.89 11.92 60.10 27.98 17.10 17.10 65.80 21.24 9.84 68.91 20.73 10.88 68.39 39.90 19.19 40.91
Qwen3-1.7b 14.51 22.28 63.21 1.55 87.56 10.88 9.33 46.63 44.04 9.33 21.76 68.91 10.36 20.21 69.43 20.20 31.82 47.98
Claude-3.7-sonnet 39.90 5.18 54.92 21.24 39.90 38.86 29.53 12.95 57.51 35.23 5.18 59.59 33.68 5.18 61.14 48.99 10.10 40.91
DeepSeek-V3 39.38 6.74 53.89 23.83 34.20 41.97 25.91 12.95 61.14 32.64 5.70 61.66 34.20 6.22 59.59 50.51 9.09 40.40
I.3 Detailed Performance by Difficulty Levels
This subsection provides the detailed average Pass@1 performance across evaluated models for SC tasks in different languages (Python, Go, TS, JS) and WSC tasks in Python, broken down by difficulty levels (Easy, Medium, Hard). The average Pass@1 is calculated as the sum of the accuracy values for all models within that difficulty level and language/task type, divided by the number of evaluated models (16 in total).
Table 26: Average Pass@1 Performance by Difficulty Level across Languages and Task Types (%).
Language/Task Type Easy (%) Medium (%) Hard (%)
SC Python 41.94 28.42 28.03
SC Go 19.75 16.50 6.40
SC TS 28.99 27.83 21.05
SC JS 30.43 28.34 20.72
WSC Python 65.75 36.11 15.12
The data presented here correspond to the aggregated results used for calculating the average Pass@1 per difficulty level shown in Table 4 in the main paper. The total number of task instances and correct completions contributing to these percentages are also shown in the raw data (provided alongside the benchmark release). This detailed breakdown allows for a finer-grained analysis of how model performance varies with the inherent complexity of the tasks within each language and task type category.
I.4 Detailed Pass Rate $\geq$ 98% Distribution
Table 27: Percentage of Problems with Test Case Pass Rate $\geq$ 98% for Evaluated Models across Task Types (%).
Model Name SC Python SC Go SC Java SC TS SC JS WSC Python
Claude-3.7-sonnet 13.47 14.69 15.58 14.18 11.40 10.10
Doubao-1-5-pro-32k-250115 12.44 13.57 14.94 9.84 11.40 9.09
Gemini-2.5-flash-preview 11.92 8.52 9.33 10.03 10.88 8.08
DeepSeek-V3 11.92 9.51 16.47 10.88 12.44 9.09
Llama-4-maverick 5.70 8.38 13.75 6.74 7.18 8.58
Qwen3-235b-a22b 12.95 10.79 10.39 10.36 12.44 9.60
Llama-4-scout 13.99 11.52 12.87 7.77 10.09 11.11
Llama-3.3-70b-Instruct 11.40 7.15 12.33 8.81 9.84 8.08
Qwen3-32b 9.84 8.56 12.64 9.33 11.92 11.62
Qwen2.5-Coder-32b-Instruct 8.81 7.65 13.79 9.84 8.81 14.65
Gemma-3-27b-it 8.81 7.72 9.95 10.88 9.33 8.08
Mistral-small-3.1-24b-Instruct 5.70 7.91 10.59 8.29 8.81 11.62
Qwen2.5-Coder-14b-Instruct 10.36 8.73 11.49 10.02 11.40 16.16
Qwen3-8b 9.84 9.91 11.34 10.88 10.88 19.19
Llama-3.2-3b-Instruct 9.54 4.17 3.59 6.86 4.93 23.74
Qwen3-1.7b 8.29 5.51 6.19 6.84 7.25 31.82
This subsection provides the detailed percentage of problems where evaluated models achieved a test case pass rate of 98% or higher for all evaluated models across all task types (SC Python, SC Go, SC Java, SC TS, SC JS, and WSC Python). This metric serves as an indicator of how frequently models produce code that is nearly correct but fails on a small number of subtle test cases generated by Property-Based Testing.
As demonstrated in Table 27, the percentage of problems where models achieved a pass rate of 98% or higher varies significantly across models and task types, generally ranging from less than 1% to over 30%. These results underscore the discriminative power of the rigorous PBT-generated test suites in identifying subtle errors that models frequently make, even when the generated code appears mostly correct.
Appendix J Limitations
Despite its strengths in generating dynamic, rigorously tested, and realistic tasks focusing on functional correctness, Code2Bench-2505, like many benchmarks, has limitations in its evaluation scope. Our primary focus is on assessing the functional correctness of generated code, measured through Pass@1 against comprehensive PBT-generated test suites. While functional correctness is paramount, real-world software development necessitates evaluating other crucial aspects of code quality, such as efficiency, readability, style, security, robustness to invalid inputs, and the ability to generate accompanying documentation or tests. Code2Bench-2505 currently does not directly evaluate these important dimensions.
Furthermore, the current iteration of Code2Bench-2505 primarily focuses on code generation tasks, where models are required to generate a complete function implementation based on a natural language instruction and function signature. However, the underlying structure of the benchmark, including the availability of ground truth implementations and the rigorous, diverse test cases generated via PBT, offers significant potential for evaluating LLM capabilities beyond simple generation. By leveraging the ground truth and PBT-generated test suites, the framework could be extended to support other task types crucial for software development workflows, such as code completion (filling in missing parts of code), code editing/repair (modifying existing code to meet new requirements or fix bugs), and assessing code reasoning abilities through execution prediction or debugging tasks. Expanding to these diverse task types would provide a more comprehensive evaluation of LLMs’ understanding and manipulation of code, moving beyond pure synthesis.
Future work could explore extending the benchmark to incorporate metrics and testing methodologies for some of these additional code quality attributes and diverse task types, providing a more holistic assessment of LLM capabilities in a full software development context.